OpenStudy (anonymous):

5log(base of)3(x)-2=8

OpenStudy (anonymous):

\[5\log _{3}(x)-2=8\]

OpenStudy (anonymous):

Thats what it should look like

OpenStudy (mathmale):

\[5\log_{3}x-2=8 \]...yes. What's the first thing you'd do to simplify this equation?

OpenStudy (anonymous):

What I did was, \[(3^{5)^{8}}=x-2\]

OpenStudy (anonymous):

I'm not sure if that is correct

OpenStudy (anonymous):

And after that, I'm stuck

OpenStudy (mathmale):

You have:\[5\log_{3}x-2=8\]... Do you mean 5(log to the base 3 of) x - 2 = 8 or do you mean 5(log to the base 3 of) (x-2)?

OpenStudy (anonymous):

I'm not sure of what you are asking. The question is 5log(base of 3)(x)-2=8

OpenStudy (mathmale):

All right. The reason I've asked that question is that your expression has -2=8 on the right, and this can be simplified by adding 2 to both sides. Do that, please. Write out the whole equation, please.

OpenStudy (anonymous):

\[5\log _{3}x=11\] Like so?

OpenStudy (mathmale):

You have\[5\log _{3}x=11\]

OpenStudy (mathmale):

Mind explaining how you got that 11?

OpenStudy (anonymous):

Yes. I did as you said with the -2 = 8 and added the 2 on each side

OpenStudy (anonymous):

I then end up with. \[(3^{5)^{11}}=x\]

OpenStudy (anonymous):

When I rearrange

OpenStudy (anonymous):

Any input?

OpenStudy (mathmale):

Don't be offended, but again I ask you to check your arithmetic. Adding 2 to both sides of your original equation results in what?

OpenStudy (anonymous):

Oh my. Right as you said that I realised my 8 is not a 9. So I have been adding 2 to 9 not 2 to 8. So yes the correct equation would be \[5\log _{3}x=10 --- (3^{5)^{10}}=x\]

OpenStudy (mathmale):

\[5\log _{3}x=10~ can be reduced ... think about how you'd do that. Leave the \log alone for now\]

OpenStudy (mathmale):

5 to the base 3 of x equals 10. Reduce. How?

OpenStudy (anonymous):

I'm not quite sure. Would the 5 go in to the 10 twice? to equal Base 3 of x equals 2

OpenStudy (mathmale):

You're making the problem unnecessarily difficult. Divide the last equation by 5, both sides. What do you get?

OpenStudy (anonymous):

cancels out on one side and 2 on the other

OpenStudy (mathmale):

And your result?

OpenStudy (anonymous):

\[3^{x}=2\]

OpenStudy (mathmale):

...\[5\log _{3}x=10\]

OpenStudy (mathmale):

Please try again: divide both sides by 2. Please don't skip steps in this procedure.

OpenStudy (anonymous):

Why am I diving both sides by 2 where did the 2 come from

OpenStudy (mathmale):

Sorry, mea culpa. Divide both sides of this equation by 5, to isolate the logarithmic expression.

OpenStudy (anonymous):

10 divided by 5 is 2, there fore i get \[\log _{3}x=2\]

OpenStudy (mathmale):

Now that's it. Right. This is a logarithmic equation. What is the base? How could you convert this to an exponential function / equation?

OpenStudy (anonymous):

Okay, thank you I got the end result of 9.

OpenStudy (mathmale):

Happy to help. You might want to go back and check your result, 9, into the original eq'n, to be certain that it's correct.

OpenStudy (anonymous):

Okay thanks :)